OpenStudy (konradzuse):

u question for the integral of 1/sqrt(9-16x^2) dx

OpenStudy (konradzuse):

I got u = 8x du = 8dx 1/8du = dx \[1/8 \int\limits \frac{du}{\sqrt{a^2 -u^2}}\]????

OpenStudy (konradzuse):

1/4*

OpenStudy (konradzuse):

hmm...

OpenStudy (konradzuse):

I actually just looked it up on wolfman, seems like I was doing it correctly. Why do I need to complete the square? http://www.wolframalpha.com/input/?i=integral+of+1%2Fsqrt%289-16x%5E2%29+dx

OpenStudy (turingtest):

you don't, that's why I deleted my remark

OpenStudy (konradzuse):

oh :P I was mainly concerned with the du, but I looked it up in the book and it seemd to be the same. I was confused at first about getting the du by itself, but I realized I can send it over to the left of the integral :).

OpenStudy (turingtest):

but u should be 4x

OpenStudy (konradzuse):

yeah I messed up :p

OpenStudy (turingtest):

besides that it's good though :)